Hello can you help me with this questions please.
Suppose that the width of a rectangle is 5 inches shorter than the lenght and that the perimeter of the rectangle is 50.
a) Set up an equation of the perimeter involving only L, the length of the rectangle.
Answer:
b) Solve this equation algebraically to find the length of the rectangle. Find the width as well.
Answer: Length ______, Width ______
Show your work here:

Answer by mayank(15)

here let width be W and length be L
From given details width is 5 inches shorter than length, thus W=L-5
now perimeter of rectangle is 50.
thus 50=2(L+W)
thus P=2(L+L-5)
thus P=2(2L-5)
thus P=4L-10
b)now as p=50, 50=4L-10;
thus 60=4L
thus L=60/4
thus L=15